3 (Chapter 6) Turlock Meats, Inc. is looking at a new processing system with an installed cost of $600,000. This cost will be depreciated straight-line to zero over the projects five-year life, at the end of which the new system can be scrapped for $100,000. The new processing system will save the firm $150,000 per year in pretax operating costs, and the system requires an initial investment in net working capital of $30,000 (which must be maintained for the five years). If the tax rate is 35% and the discount rate is 10%, what is the NPV of the project?
Answer #1

Initial Investment=600000

Initial Working Capital Investment=30000..This will be recovered at the end of year 5

Free Cash flows=(150000-600000/5)*(1-35%)+600000/5=139500

After-tax salvage value=100000*(1-35%)=65000

So, NPV=-600000-30000+139500/1.1+139500/1.1^2+139500/1.1^3+139500/1.1^4+139500/1.1^5+30000/1.1^5+65000/1.1^5=-42197.72